The Jurassic Zoo charges ​$6 for each adult admission and ​$4 for each child. The total bill for the 101 people from a school trip was ​$444. How many adults and how many children went to the​ zoo?  
Answer by Boreal(15235)

A+c=101, count equation
6A+4c=444, money equation
-4A-4c=-404
2A=40
A=20, so c=81
20 adults ($120)
81 children ($324)
adds to $444